OpenStudy (anonymous):

I need to find the square root radical so the problem is square root of 8x - 7 then -7=0

OpenStudy (anonymous):

\[\sqrt{8x-7}-7=0\]

OpenStudy (campbell_st):

well rewrite the equation \[\sqrt{8x - 7}=7\] square both sides \[8x -7 = 49\] add 7 to both sides \[8x = 56\] hopefully this helps